# Wildfire Risk Prediction: A Review

### v3

@@ 6 Fire danger rating systems
For example, Canada developed a meteorology-based fire danger rating method in 1968, known as the Canadian Forest Fire Danger Rating System (CFFDRS). This system comprises four main subsystems: the Fire Weather Index (FWI), Fire Behavior Prediction (FBP), Fire Occurrence Prediction (FOP), and Auxiliary Fuel Moisture (AFM) (Van Wagner et al., 1974; Stocks et al., 1989b; Lawson and Armitage, 2008). Among these, the FWI system is the most widely used (Fig. 4). The FWI system, based on the theory of time lag and equilibrium moisture content, calculates changes in fuel moisture codes (i.e., Fine Fuel Moisture Code (FFMC), Duff Moisture Code (DMC), and Drought Code (DC)) and fire behavior indices (i.e., Initial Spread Index (ISI) and Buildup Index (BI)) according to weather conditions. These indices are then used to classify potential wildfire danger levels in forests based on the moisture content of fuels at different locations or sizes (Stocks et al., 1989a).
@@
In comparison, the United States National Fire Danger Rating System (USNFDRS) (Deeming, 1972) evaluates fire risk by considering not only meteorological factors but also various terrain types, fuel types, and fuel models. By inputting meteorological data, fuel moisture, and terrain into the model, four output components are obtained: the Spread Component (SC), Energy Release Component (ERC), Burning Index (BI), and Ignition Component (IC). The first three indices, based on combustion physics, correspond to fire behavior characteristics, while the IC provides an estimate of fire danger levels (Zacharakis and Tsihrintzis, 2023a). The system was revised in 1978 and 1988 to address issues such as poor response to prolonged droughts and the inability to capture real-time fuel dynamics (Deeming et al., 1977) and inadequate performance in moist environments (Burgan, 1988). Recently, Jolly et al. (2024) updated the USNFDRS introduced in 1988 by emphasizing the provision of more detailed, fuel-type-dependent combustion condition assessments, proposing a new fuel moisture model, and simplifying the fuel model.
@@ 7 Deep Learning Methods for Wildfire Risk Prediction
Currently, popular methods for predicting wildfire risk include fire rating systems, statistical methods, traditional machine learning methods, and deep learning approaches. Fire rating systems, developed and extensively applied in regions like Canada, the USA, and Australia, are based on a combination of empirical and mathematical modeling principles involving fluid dynamics, combustion, energy transfer, and considerations of fuel, weather, and terrain conditions (Rezaie et al., 2023; Mell et al., 2007). These models are known for their interpretability and generalizability, particularly suitable when training data are limited. However, physical models heavily rely on implicit assumptions about soil moisture, fuel load, land cover types, landscape structure, air humidity, and associated fire behaviors, which could adversely affect prediction accuracy (Rezaie et al., 2023; Bar Massada et al., 2011).
For small-scale or specific target analyses of wildfire risk, three additional methods are commonly employed. Firstly, statistical methods use various regression models to establish relationships between observed wildfire risks and factors such as fuel conditions, meteorological and climatic conditions, and terrain features. These methods are relatively simple but have limited learning capabilities and tolerance for the randomness and high nonlinearity typical of wildfire prediction tasks (Li et al., 2022a). In contrast, traditional machine learning methods like Random Forests (RF), Support Vector Machines (SVM), ensemble learning models, and Bayesian models automate the modeling of patterns and regularities in historical fire data and various wildfire risk factors through data-driven approaches and manual feature engineering. They handle non-linearities with higher interpretability and significantly better predictive capabilities than statistical methods. However, many of these models lack good interpretability.
Over the past decade, wildfire modeling has evolved from using statistical algorithms to more advanced machine learning methods that can automatically improve through experience (Jain et al., 2020; Oliveira et al., 2021b). The advantages of statistical methods include simplicity in modeling, low data requirements, and strong interpretability. However, limited by a small number of parameters, statistical methods often have poor predictive performance and generalization capabilities. Machine learning methods are increasingly used due to their ability to identify distributions in large datasets without the need for prior knowledge of the phenomena and to model in a non-parametric and non-linear manner (Oliveira et al., 2021b). Yet, their high dimensionality and complexity pose challenges in interpreting the importance of independent variables compared to traditional statistical regression methods. According to the classification proposed by Jain et al. (2020), common machine learning models used for wildfire prediction include decision trees, SVM, ANN, Bayesian methods, reward-based methods, clustering methods, and others. Some researchers have found that decision tree methods, such as Random Forests, perform better in wildfire prediction than other models. For instance, Malik et al. (2021) observed that Random Forest algorithms outperformed all other tested algorithms in terms of accuracy across various target labels and dataset subsets. However, some researchers argue that different methods have their advantages and disadvantages, and it is not conclusive that a single model is superior to others (Jaafari et al., 2019; Razavi Termeh et al., 2018). Therefore, some studies combine predictions from two or more models to improve final prediction accuracy. Since Jain et al. (2020); Zacharakis and Tsihrintzis (2023b) have already extensively reviewed the application of statistical methods and machine learning algorithms in wildfire prediction, detection, fuel characteristics, and other aspects of wildfire science and management up to the years 2020 and 2023, this section will not repeat those details.
The traditional machine learning methods mentioned earlier typically use shallow architectures and fail to fully leverage the spatial patterns inherent in images (Zhang et al., 2018). These classifiers directly categorize input data, lacking an effective feature extraction process, which limits their ability to extract representative features and improve classification accuracy (Zhang et al., 2019). Compared to traditional machine learning, deep learning models, although requiring more training data, can automatically extract more discriminative predictive features and learn more complex relationships and distributions through hierarchical structures and activation functions. This reduces reliance on manual feature engineering and enhances generalizability, making them better suited for complex real-world conditions. Additionally, deep learning algorithms have a larger parameter capacity, enabling them to handle larger datasets and reducing the risk of model overfitting. While deep learning models have become an increasingly popular tool for predicting wildfire risks and represent a new trend in the field, current literature reviews, such as those by Jain et al. (2020); Zacharakis and Tsihrintzis (2023b), provide only a limited overview of methods including Recurrent Neural Network (RNN) and CNN, lacking a comprehensive systematization of their use.
To address this gap, this section systematically reviews the latest advancements in deep learning-based wildfire risk prediction models. Currently, wildfire risk prediction using deep learning methods generally transforms the task into segmentation or classification challenges, where models are trained using historical wildfire data and driving factors affecting fire ignition or spread as inputs, with future wildfire occurrences or risk levels as ground truth. For segmentation tasks, models can be categorized based on how they use the data into three types: time series prediction, classic semantic segmentation, and spatio-temporal prediction. Time series prediction uses 1D feature vectors of time series as input and employs recurrent neural network methods like RNN, LSTM, and GRU, as well as attention-based Transformer methods, to predict wildfire risks at the next time step. Classic semantic segmentation methods take 2D feature maps of wildfire influencing factors or wildfire data from a previous time step as inputs to predict the next time step’s wildfire risks, outputting results in 2D. These models are typically CNNs or Transformers. Spatio-temporal prediction methods combine the above approaches, using 2D time series data of wildfire influencing factors or wildfire data as inputs to predict the next time step’s wildfire risks. Deep learning models used for spatio-temporal prediction often combine models such as CNNs, Transformers, GCNs, and GNNs with RNNs, LSTM, and GRU in various ways. Thus, this section will introduce models based on time series prediction methods (including RNN, LSTM, GRU, and Transformers), classic semantic segmentation methods (CNNs), and spatio-temporal prediction methods for wildfire risk. Furthermore, since deep learning models are often considered black box models, making it difficult to interpret the contribution of various features to model predictions and the prediction process itself, this section also establishes techniques for explainability of deep learning models to support decision-making in wildfire management and model development.
@@ 7.1 Wildfire risk rating and probability calibration
Even though Jain et al. (2020); Zacharakis and Tsihrintzis (2023b) systematically reviewed the application of various machine learning methods in wildfire risk research, but they did not provide a description of the post-processing techniques required to convert the predictions of machine learning models into actual wildfire risk. Directly using the output of these models as the probability of wildfire risk may lack theoretical foundation and stability. This is because the machine learning model, trained using the wildfire coverage at time $t+n$ as the ground truth and input data from time $t$ such as fuel and environmental conditions, inherently outputs a probability map that reflects the likelihood of a specific location being within the wildfire area at time $t+n$, rather than representing actual risk. Additionally, even with identical inputs, different models may produce significantly divergent probability maps.
@@
Moreover, wildfire historical data sets typically suffer from severe class imbalance, where the spatial and temporal scope of fire occurrences is considerably smaller compared to non-fire events. To expedite model convergence, many machine learning datasets are often constructed with a balanced class ratio, typically 1:1, 1:1.5, or 1:2 (Huot et al., 2021; Kondylatos et al., 2022) between fire and non-fire instances. This may lead to an overestimation of false positives during model evaluation (Bakke et al., 2023). However, actual wildfire risk probabilities are usually much lower. For example, in wildfire risk assessments for the Mediterranean region and the United States, Ager et al. (2014a) and Preisler et al. (2004) respectively, used probabilities in the thousandths range.
@@
Using the probability output of machine learning models as wildfire risk could be misleading to non-experts. Therefore, many studies convert continuous wildfire risk probabilities into risk levels to support wildfire management, such as conveying public warnings and guiding fire suppression measures. Among the methods for risk level classification, the Jenks natural breaks classification (Jenks and Caspall, 1971) is most commonly used. For instance, Xie et al. (2022b) trained an ensemble machine learning model using binary wildfire ignition records and then classified the predicted probabilities into five groups using the Jenks method. Similarly, Iban and Sekertekin (2022) and Moayedi and Khasmakhi (2023) also classified the hazard maps generated by machine learning models into five categories using the Jenks method. Along the same lines, Chen et al. (2024c) used the Jenks method to divide the estimated potential rate of spread and fire radiative power into five levels for wildfire risk assessment. Other methods have also been employed for risk level classification. For example, Liang et al. (2019) used the Kennard-Stone method combined with standardized duration and fire size to categorize the wildfire risk predicted by an RNN into five levels. Jalilian and Jouibary (2023) applied natural discontinuity classification methods to group the probabilities predicted by machine learning models into three classes. Bjånes et al. (2021) and Chicas et al. (2022) divided the wildfire ignition probability maps produced by the models into five levels using an equal-interval method, with intervals of 0.2. Similarly, Trucchia et al. (2022) manually defined five risk levels.
@@
From the above analysis, it is evident that the natural breaks method is commonly used to classify machine learning-predicted probability maps into multiple levels. However, there is still limited exploration on how to map these probability maps into continuous probability distributions and establish a relationship between predicted probabilities and actual risk probabilities.
@@ 7.2 Time series prediction
7.2.1 Recurrent neural network
As discussed in previous sections, the occurrence and spread of wildfires are influenced by various factors, including meteorological conditions, fuel status, and historical wildfire data. These factors exhibit cumulative effects. Therefore, utilizing time series data enables models to capture the gradual changes in influencing factors, thus facilitating the prediction of future wildfire risks. The most classical methods in time series prediction tasks are Recurrent Neural Networks (RNNs), which include RNN, LSTM, and Gated Recurrent Unit (GRU).
@@ 7.2.1 Recurrent neural network
In wildfire risk prediction, classic RNN networks and their variants have been employed. For instance, Elman RNN designed a spatiotemporal prediction framework integrating historical wildfire and weather data to predict the annual average burned area in Canadian forests. Experiments demonstrated a prediction error of less than 0.5ha on areas of 500ha, validating the feasibility of using RNN for fire prediction. Similarly, a comparative study of ten artificial neural networks (ANNs) based on meteorological data for predicting wildfire incidence in Australia found that Elman RNN exhibited the best performance, with average accuracy, sensitivity, and specificity exceeding 93%. Further, a network composed of a dynamic autoencoder and RNN was used to predict the next month’s burned area in five U.S. regions, where the dynamic autoencoder was employed to transform multidimensional time series data into RNN input features, and dual decoders were used for predicting fire spot likelihood over different time spans, i.e., one week and one month. Comparative experiments showed that their method was slightly superior to comparative methods, generative network, and GRU.
While Jain et al. (2020) and others have highlighted the limited use of classical RNNs, our findings are consistent with these observations. The underutilization of RNNs may be attributed to their difficulties in handling long sequence data, where they are prone to gradient vanishing or exploding, making it challenging to capture long-term dependencies (Hochreiter, 1997; Gers et al., 2000). To address these issues, LSTM networks were developed. As illustrated in Fig. 7, LSTMs incorporate three gating units: input gate, forget gate, and output gate, which control the flow and retention of information. These gates, with trainable weights, decide whether to pass, forget, or output information based on the current input and the output from the previous timestep. This gating mechanism effectively resolves the issues of vanishing and exploding gradients, thereby enabling better handling of long sequences and capturing of long-term dependencies. Consequently, LSTMs are more prevalently applied in wildfire risk prediction than classical RNNs. The specific configuration of an LSTM cell is depicted in Fig. 7b and Eq.14 (Monner and Reggia, 2012; Liang et al., 2019).
@@
Studies utilizing LSTM for wildfire risk prediction often model various risk factors, including weather and climate conditions, fuel conditions, and socio-economic factors. For instance, Liang et al. (2019) found that LSTM outperformed RNN and Back Propagation Neural Network (BPNN) in predicting the class of wildfires in Alberta, Canada, based on meteorological conditions, achieving an overall accuracy of 90.9%. Similar findings were reported by Natekar et al. (2021), who assessed the wildfire risk in Indian forests using meteorological parameters and LSTM methods. The accuracy and RMSE were 94.77% and 37.5%, respectively, surpassing other benchmark methods like CNN and SVM. Additionally, Kondylatos et al. (2022) compared RF, XGBoost, and deep learning methods (LSTM and ConvLSTM) and found that deep learning significantly outperformed traditional machine learning methods in predicting next day wildfires in the Mediterranean using ERA5-Land meteorological data, MODIS NDVI, and diurnal LST data along with soil moisture, terrain data, land cover types, and distances from infrastructure. Li et al. (2023a) further integrated attention mechanisms and deep learning interpretability into an LSTM network, using monthly-scale climate, socio-economic, fuel variables, and oceanic indices to predict short-term (1–4 months) and long-term (5–8 months) wildfire burning areas in tropical regions. Experimental results demonstrated that the enhanced LSTM model outperformed traditional machine learning methods and the baseline LSTM model. Additionally, interpretability analysis revealed that precipitation and vapor pressure deficit were critical driving factors, while oceanic indices provided more significant contributions to long-term wildfire burning area predictions.

### v4

@@ 6 Fire danger rating systems
For example, Canada developed a meteorology-based fire danger rating method in 1968, known as the Canadian Forest Fire Danger Rating System (CFFDRS). This system comprises four main subsystems: the FWI, Fire Behavior Prediction (FBP), Fire Occurrence Prediction (FOP), and Auxiliary Fuel Moisture (AFM) (Van Wagner et al., 1974; Stocks et al., 1989b; Lawson and Armitage, 2008). Among these, the FWI system is the most widely used (Fig. 4). The FWI system, based on the theory of time lag and equilibrium moisture content, calculates changes in fuel moisture codes (i.e., Fine Fuel Moisture Code (FFMC), Duff Moisture Code (DMC), and Drought Code (DC)) and fire behavior indices (i.e., Initial Spread Index (ISI) and Buildup Index (BI)) according to weather conditions. These indices are then used to classify potential wildfire danger levels in forests based on the moisture content of fuels at different locations or sizes (Stocks et al., 1989a).
@@
In comparison, the United States National Fire Danger Rating System (USNFDRS) (Deeming, 1972) evaluates fire risk by considering not only meteorological factors, but also various terrain types, fuel types, and fuel models. By inputting meteorological data, fuel moisture, and terrain into the model, four output components are obtained: the Spread Component (SC), Energy Release Component (ERC), Burning Index, and Ignition Component (IC). The first three indices, based on combustion physics, correspond to fire behavior characteristics, while the IC provides an estimate of fire danger levels (Zacharakis and Tsihrintzis, 2023a). The system was revised in 1978 and 1988 to address issues such as poor response to prolonged droughts and the inability to capture real-time fuel dynamics (Deeming et al., 1977) and inadequate performance in moist environments (Burgan, 1988). Recently, Jolly et al. (2024) updated the USNFDRS introduced in 1988 by emphasizing the provision of more detailed, fuel-type-dependent combustion condition assessments, proposing a new fuel moisture model, and simplifying the fuel model.
@@ 7 Deep Learning Methods for Wildfire Risk Prediction
Currently, methods for wildfire risk prediction include fire rating systems, statistical models, traditional machine learning, and deep learning approaches. Fire rating systems, extensively applied in regions like Canada, the USA, and Australia, combine empirical and mathematical modeling based on fluid dynamics, combustion, energy transfer, and factors like fuel, weather, and terrain conditions (Rezaie et al., 2023; Mell et al., 2007). These models are interpretable and generalizable, especially when training data is limited. However, their reliance on assumptions regarding soil moisture, fuel load, landscape structure, and other factors can reduce prediction accuracy (Bar Massada et al., 2011).
For more specific analyses, statistical models use regression to establish relationships between wildfire risks and influencing factors such as fuel conditions, meteorology, and terrain. While simple and interpretable, they have limited capacity for handling the nonlinearity and randomness typical in wildfire prediction tasks (Li et al., 2022a). Traditional machine learning methods (e.g., Random Forests, SVM, ensemble models, and Bayesian methods) improve upon statistical models by using data-driven techniques and manual feature engineering to identify patterns in historical fire data. These methods operate non-parametrically and are better suited for non-linearities (Oliveira et al., 2021b), yet they present challenges in variable interpretation compared to statistical approaches. Over the past decade, the focus in wildfire modeling has shifted from statistical algorithms to machine learning (Jain et al., 2020; Oliveira et al., 2021b). Decision tree methods, such as Random Forests (RF), are particularly popular due to their accuracy and interpretability, as highlighted in studies like Malik et al. (2021). Nevertheless, some researchers argue that no single model is universally superior, suggesting a need for model combinations to enhance accuracy (Jaafari et al., 2019; Razavi Termeh et al., 2018).
Traditional machine learning methods often use shallow architectures that fail to exploit spatial patterns effectively, limiting feature extraction and classification accuracy (Zhang et al., 2018). In contrast, deep learning models, despite their higher data requirements, can automatically extract complex features through hierarchical structures, reducing the need for manual feature engineering and enhancing generalizability. Their larger parameter capacity also allows them to manage large datasets and minimize overfitting risks. Although deep learning has become increasingly popular for wildfire risk prediction, reviews such as Jain et al. (2020) and Zacharakis and Tsihrintzis (2023b) have only partially covered models like Recurrent Neural Networks (RNNs) and CNNs without a comprehensive systematization.
This section reviews the latest advancements in deep learning-based wildfire risk prediction models, which are typically framed as segmentation or classification tasks. These models use historical wildfire data and influencing factors as inputs, with future wildfire occurrences or risk levels as ground truth. Models are categorized into time series prediction, image semantic segmentation and classification, and spatio-temporal prediction. Time series models (RNN, LSTM, Gated recurrent unit (GRU), Transformers) predict wildfire risks using 1D feature vectors, while semantic segmentation and classification models (CNNs, Transformers) employ 2D feature maps to produce 2D risk predictions. Spatio-temporal models integrate these approaches using 2D time series data.
Deep learning models for spatio-temporal prediction often combine CNNs, Transformers, GCNs, and graph neural networks (GNNs) with RNNs, LSTMs, or GRUs, enhancing resolution with higher-resolution data over extended sequences. Despite advances in GPU technology supporting larger data scales, these models still struggle with efficiency and large spatiotemporal scales compared to newer models like Mamba (Gu and Dao, 2024). Additionally, many models use multi-branch networks that rely solely on images or nodes, missing the benefits of multimodal approaches like incorporating textual information.
While previous reviews (e.g., Jain et al. (2020); Zacharakis and Tsihrintzis (2023b)) have examined machine learning in wildfire science, they often omit post-processing techniques essential for converting model outputs into actionable risk information. Directly using these outputs can be misleading due to class imbalances and variability between models. This section also introduces methods for classifying and calibrating deep learning wildfire risk predictions to address these challenges.
@@ 7.1 Wildfire risk rating and probability calibration
Although Jain et al. (2020); Zacharakis and Tsihrintzis (2023b) systematically reviewed the application of various machine learning methods in wildfire risk research, they did not provide a description of the post-processing techniques required to convert the predictions of machine learning models into actual wildfire risk. Directly using the output of these models as the probability of wildfire risk may lack theoretical foundation and stability. This is because wildfire historical data sets typically suffer from severe class imbalance, where the spatial and temporal scope of fire occurrences is considerably smaller compared to non-fire events. To expedite model convergence, many machine learning datasets are often constructed with a balanced class ratio, typically 1:1, 1:1.5, or 1:2 (Huot et al., 2021; Kondylatos et al., 2022) between fire and non-fire instances. This may lead to an overestimation of false positives during model evaluation (Bakke et al., 2023). However, actual wildfire risk probabilities are usually much lower. For example, in wildfire risk assessments for the Mediterranean region and the United States, Ager et al. (2014a) and Preisler et al. (2004) used probabilities in the thousandths range, respectively.
@@
Using the probability output of machine learning models as wildfire risk could be misleading to non-experts. Therefore, many studies convert continuous wildfire risk probabilities into risk levels to support wildfire management, such as conveying public warnings and guiding fire suppression measures. Among the methods for risk level classification, the Jenks natural breaks classification (Jenks and Caspall, 1971) is most commonly used. For instance, Xie et al. (2022) trained an ensemble machine learning model using binary wildfire ignition records and then classified the predicted probabilities into five groups using the Jenks method. Similarly, Iban and Sekertekin (2022) and Moayedi and Khasmakhi (2023) also classified the hazard maps generated by machine learning models into five categories using the Jenks method. Along the same lines, Chen et al. (2024c) used the Jenks method to divide the estimated potential rate of spread and fire radiative power into five levels for wildfire risk assessment.
Other methods have also been employed for risk level classification. For example, Liang et al. (2019) used the Kennard-Stone method combined with standardized duration and fire size to categorize the wildfire risk predicted by an RNN into five levels. Jalilian and Jouibary (2023) applied natural discontinuity classification methods to group the probabilities predicted by machine learning models into three classes. Bjånes et al. (2021) and Chicas et al. (2022) divided the wildfire ignition probability maps produced by the models into five levels using an equal-interval method, with intervals of 0.2. Similarly, Trucchia et al. (2022) manually defined five risk levels.
@@ 7.2 Time Series Prediction
7.2.1 Recurrent Neural Networks
As discussed in previous sections, the occurrence and spread of wildfires are influenced by various factors, including meteorological conditions, fuel status, and historical wildfire data. These factors exhibit cumulative effects. Therefore, utilizing time series data enables models to capture the gradual changes in influencing factors, thus facilitating the prediction of future wildfire risks. The most classical methods in time series prediction tasks are RNNs, which include RNN, LSTM, and GRU.
@@ 7.2.1 Recurrent Neural Networks
Classic RNN networks and their variants have been employed in wildfire risk prediction. For instance, Cheng and Wang (2008) designed a spatiotemporal Elman RNN-based prediction framework integrating historical wildfire and weather data to predict the annual average burned area in Canadian forests. Experiments demonstrated a prediction error of less than 0.5ha on areas of 500ha, validating the feasibility of using RNN for fire prediction. Similarly, a comparative study of ten artificial neural networks (ANNs) based on meteorological data for predicting wildfire incidence in Australia found that Elman RNN exhibited the best performance, with average accuracy, sensitivity, and specificity exceeding 93% (Dutta et al., 2013). Furthermore, a network composed of a dynamic autoencoder and RNN was used to predict the next month’s burned area in five U.S. regions, where the dynamic autoencoder was employed to transform multidimensional time series data into RNN input features, and dual decoders were used for predicting fire spot likelihood over different time spans, i.e., one week and one month. Comparative experiments showed that their method was slightly superior to comparative methods, generative network, and GRU (Chavalithumrong et al., 2021).
While Jain et al. (2020) and others have highlighted the limited use of classical RNNs, our findings are consistent with these observations. The underutilization of RNNs may be attributed to their difficulties in handling long sequence data, where they are prone to gradient vanishing or exploding, making it challenging to capture long-term dependencies (Hochreiter, 1997; Gers et al., 2000). To address these issues, LSTM networks were developed. As illustrated in Fig. 7, LSTMs incorporate three gating units: input gate, forget gate, and output gate, which control the flow and retention of information. These gates have trainable weights and decide whether to pass, forget, or output information based on the current input and the output from the previous time steps. This gating mechanism effectively resolves the issues of vanishing and exploding gradients, thereby enabling better handling of long sequences and capturing of long-term dependencies. Consequently, LSTMs are more prevalently applied in wildfire risk prediction than classical RNNs. The specific configuration of an LSTM cell is depicted in Fig. 7b and Eq.13 (Monner and Reggia, 2012; Liang et al., 2019).
@@
Studies utilizing LSTM for wildfire risk prediction often model various risk factors, including weather and climate conditions, fuel conditions, and socio-economic factors. For instance, Liang et al. (2019) found that LSTM outperformed RNN and Back Propagation Neural Network (BPNN) in predicting the class of wildfires in Alberta, Canada, based on meteorological conditions, achieving an overall accuracy of 90.9%. Similar findings were reported by Natekar et al. (2021), who assessed the wildfire risk in Indian forests using meteorological parameters and LSTM methods. The accuracy and RMSE were 94.77% and 37.5%, respectively, surpassing other benchmark methods like CNN and SVM. Additionally, Kondylatos et al. (2022) compared RF, XGBoost, and deep learning methods (LSTM and ConvLSTM) and found that deep learning significantly outperformed traditional machine learning methods in predicting next day wildfires in the Mediterranean using ERA5-Land meteorological data, MODIS NDVI, and diurnal LST data along with soil moisture, terrain data, land cover types, and distances from infrastructure. Li et al. (2023b) further integrated attention mechanisms and deep learning interpretability into an LSTM network, using monthly-scale climate, socio-economic, fuel variables, and oceanic indices to predict short-term (1–4 months) and long-term (5–8 months) wildfire burning areas in tropical regions. Experimental results demonstrated that the enhanced LSTM model outperformed traditional machine learning methods and the baseline LSTM model. Furthermore, Cheng et al. (2022, 2024) have combined LSTM with reduced order modeling techniques to improve the prediction accuracy for both regional and global scale wildfire systems. Additionally, interpretability analysis revealed that precipitation and vapor pressure deficit were critical driving factors, while oceanic indices provided more significant contributions to long-term wildfire burning area predictions.